a square poster has sides measuring 2 feet less than the sides of a square sign. if the difference between their areas is 32 square feet, find the lengths of the sides of the poster and the sign
Answer by cleomenius(959)

The area of the sign minus the area of the poster is given as 32 sq ft.
The equation will be:
(x + 2)^2 - x^2 = 32
x^2 + 4x + 4 - x^2 = 32
4x + 4 = 32
4x = 28
x = 7 ft, the side of the poster.
x + 2 = 9 ft, the size of the sign
Too check:
81 sq ft minus 49 sq ft = 32 sq ft.
Cleomenius.